OpenStudy (anonymous):

PLEASE HELP I DONT UNDERSTAND AM I AM HORRIBLE AT GRAPHING....Recall the equation for projectiles is y = –16x2 + vx + s, where y is the object’s current height (in feet) at any time after it is projected, x is the time (in seconds) the object has been in the air, v is the object’s starting upward velocity, and s is the object’s starting height.

OpenStudy (anonymous):

Do you need to graph it? Or something else?

OpenStudy (anonymous):

use graphing technology to complete this activity since many of the solutions will most likely involve irrational numbers.... I REALLY DONT GET IT AT ALL. i really need help, do you get it ?

OpenStudy (anonymous):

What is one of the questions that you are working on that uses this equation?

OpenStudy (anonymous):

That may give you the correct answer, but it will make you worse off in the long run.

OpenStudy (anonymous):

yea i want to learn how to do it not just get the answer, lol i do take my eoc's and want a future !

OpenStudy (anonymous):

@aacehm im going to ask my teacher about that one, i have others would u like to help ?

OpenStudy (anonymous):

Sure, go ahead

OpenStudy (anonymous):

Motor a speed is 190 ft per sec, motor b is 210 ft per secs. and motor c is 230 ft per sec. You have various motor speed options but need to choose the one that most closely meets the client’s height request of 690 feet. useing the projctile formula y=-16x^2+vx+s w/ starting height of 0(s=0)

OpenStudy (anonymous):

Okay, so how do you think you should start?

OpenStudy (anonymous):

and then it says Substitute the data provided for speed (v) and starting heights (s) into their proper positions in the formula y = –16x2 + vx + s and write the equation corresponding with each motor in the space provided below. Then identify the a, b, and c values for each where indicated. (6 points total, 2 points each) The equation for motor A is: a = b = c = The equation for motor B is: a = b = c = The equation for motor C is: a = b = c = what do i do ?

OpenStudy (anonymous):

i assume i start with the formula they provide

OpenStudy (anonymous):

okay, for the a values, they are all the same. Do you remember the form for quadratic equations?

OpenStudy (anonymous):

ax2 + bx + c = 0 ???

OpenStudy (anonymous):

right. So according to the equation -16x^2 + vx + s, what is a?

OpenStudy (anonymous):

-16 ?

OpenStudy (anonymous):

right! Okay, so now on to c, because that's easy too. What does c correspond to in the equation -16x^2+vx+s?

OpenStudy (anonymous):

vx ?

OpenStudy (anonymous):

nonoo i mean s

OpenStudy (anonymous):

yes, s. And in the problem, the starting height is zero, so what is c for each equation?

OpenStudy (anonymous):

i think 0

OpenStudy (anonymous):

yes, good job! Okay, so now where do you think the speed of the motor goes in the equation?

OpenStudy (anonymous):

um vx?

OpenStudy (anonymous):

im not sure on that one

OpenStudy (anonymous):

well, in this case you need to keep x, so you substitute the motor speed for v. Lastly, where does the 690 go?

OpenStudy (anonymous):

after the = sign since thats the guys expetations ?

OpenStudy (anonymous):

no, the desired height is 690, so you substitute 690 for y. Now, putting everything together, what is the equation for motor a?

OpenStudy (anonymous):

-16x^2+190x+0=690 ?

OpenStudy (anonymous):

Yes! Good job! Alright, so do you know how to solve quadratic equations? How would you solve this one?

OpenStudy (anonymous):

im really not sureee....:// i feel dumb

OpenStudy (anonymous):

Okay, here we go.\[-16x^{2}+190x=690\]\[-16x^{2}+190x-690=0\]

OpenStudy (anonymous):

Now, what can you factor out of this equation. In other words, what is the GCF of 16, 190, and 690?

OpenStudy (anonymous):

2 ?

OpenStudy (anonymous):

Yeah so factoring out a -2 would be good, since it's easier when a is positive. then we get \[-2(8x^{2}-95x+345)=0\] So now we have to factor. Do you remember how to factor? What two numbers multiply to 345 and add to -95? This is kind of tricky

OpenStudy (anonymous):

actually nevermind don't do that

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

just use the quadratic formula, do you remember that?

OpenStudy (anonymous):

ax2 + bx + c = 0

OpenStudy (anonymous):

no, the one \[\frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

oh my,i dont recell how to do that one, but yes im familiar with it

OpenStudy (anonymous):

actually, you can use a graphing calculator right? graph all three of the equations\[-16x^{2}+190x-690\]\[-16x^{2}+210x-690\]\[-16x^{2}+230x-690\]

OpenStudy (anonymous):

I dont have one, and the one downloaded on my labtop isn't working for some reason. to turn my work in i have to draw it on fraph paper and scan it but im horrible with graphs

OpenStudy (anonymous):

ugh. this is pretty tough.

OpenStudy (anonymous):

you really need a graphing calculator for it

OpenStudy (anonymous):

im sorry, you dont have to help me if you dont want. but i thank you for how far youve gotten me so far. appreciate it much

OpenStudy (anonymous):

wait hold on ima see if they have a free one online

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

so if you graph the following three functions \[-16x^{2}+190x\]\[-16x^{2}+210x\]\[-16x^{2}+230x\]

OpenStudy (anonymous):

that is what the rockets' paths would look like if they were actually fired

OpenStudy (anonymous):

so which one is closest to 690 at the peak of its arc?

OpenStudy (anonymous):

it keeps saying error, can you maybe post a pic of them graphed ?

OpenStudy (anonymous):

wait im not suppost to put the -690 ?

OpenStudy (anonymous):

well, if you put the 690 you would just have to see which is closest to 0 at its peak instead of 690, whichever one you prefer

OpenStudy (anonymous):

its the motor a ?

OpenStudy (anonymous):

no, if you look the peak of motor b is actually closest

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so that's the answer, motor b

OpenStudy (anonymous):

oh alright thank you i appreiate your help !

OpenStudy (anonymous):

ONE MORE THING, how woulldi find the vertex coordinates by looking at the graphs @aacehm

OpenStudy (anonymous):

no problem um. you just do -b/2a and f(-b/2a)

OpenStudy (anonymous):

Thank you @aacehm can you help me wit one more ?